OpenStudy (anonymous):

what is the value of round((4.3+.8)/-2-2)

OpenStudy (compassionate):

5.1/-2-2 5.1/-4 =?

OpenStudy (anonymous):



OpenStudy (anonymous):

As it is 'round'

OpenStudy (anonymous):

What does it mean, round?

OpenStudy (anonymous):

\[\frac{ 4.3+0.8 }{ -2-2 } = \frac{ 5.1 }{ -4 } \]

OpenStudy (anonymous):

= 1.275 = 1.28 ( 3 s.f.)

OpenStudy (anonymous):

Which you "round" to 1?